The corona disease has been led to many deaths per week increases by 9.2% every week.Suppose this trend does not change and 82 deaths have been recorded this week.Determined the number of deaths that: a)were recorded 4 weeks ago b) will be recorded in 3 weeks time c) will be recorded in 6 weeks time





1
Expert's answer
2022-03-14T17:34:32-0400

Solution

The deaths follow an exponential growth and can be calculated using the formula


"P=P_0(1+r)^t"

where

"P" is the number of deaths after t weeks

"P_0" is the current number of deaths

"r" is the rate of increase which "=0.092"


(a) Cases 4 weeks ago

"P_{-4}=P_0(1+r)^t"

"P_{-4}=82(1-0.092)^4"

"=56" deaths


(b) Cases in 3 weeks time

"P_3=P_0(1+r)^t"

"P_3=82(1+0.092)^3"

"=107" deaths


(c) Cases in 6 weeks time

"P_6=P_0(1+r)^t"

"P_6=82(1+0.092)^6"

"=139" deaths